# Pretrain Finite Element Method: A Pretraining and Warm-start Framework for PDEs via Physics-Informed Neural Operators

### v1

@@ Abstract
We propose a Pretrained Finite Element Method (PFEM), a physics-driven framework that bridges the efficiency of neural operator learning with the accuracy and robustness of classical finite element methods (FEM). PFEM consists of a physics-informed pretraining stage and an optional fine-tuning stage. In the pretraining stage, a neural operator based on the Transolver architecture is trained solely from governing partial differential equations, without relying on labeled solution data. The model operates directly on unstructured point clouds, jointly encoding geometric information, material properties, and boundary conditions, and produces physically consistent initial solutions with extremely high computational efficiency. PDE constraints are enforced through explicit finite element–based differentiation, avoiding the overhead associated with automatic differentiation. In the fine-tuning stage, the pretrained prediction is used as an initial guess for conventional FEM solvers, preserving their accuracy, convergence guarantees, and extrapolation capability while substantially reducing the number of iterations required to reach a prescribed tolerance. PFEM is validated on a broad range of benchmark problems, including linear elasticity and nonlinear hyperelasticity with complex geometries, heterogeneous materials, and arbitrary boundary conditions. Numerical results demonstrate strong generalization in the pretraining stage with relative errors on the order of 1%, and speedups of up to one order of magnitude in the fine-tuning stage compared to FEM with zero initial guesses.
@@
Transolver , Neural operator , Physics-informed neural operator , Physics-informed neural network , AI for PDEs
@@ 5.5 Unlimited generalization and self-learning capability of PFEM
PFEM possesses an inherent capability for unlimited generalization. This property arises from the fact that, in the fine-tuning stage, PFEM employs conventional FEM-based iterative solvers, which naturally exhibit extrapolation capability beyond the training distribution. Consequently, the performance of PFEM does not fundamentally deteriorate when encountering unseen geometries, material properties, or boundary conditions. In practice, the number of iterations required in the fine-tuning stage primarily depends on the quality of the initial solution provided by the pretraining stage.
To reduce the computational cost of the fine-tuning stage, it is therefore crucial that the pretraining phase delivers a high-quality initial guess. Notably, the pretraining phase of PFEM does not require labeled data; it only relies on point-cloud inputs encoding geometry, material properties, and boundary conditions. This observation highlights an important future research direction: how to optimally sample the input space during pretraining?
@@
In summary, the higher the accuracy achieved in the pretraining phase of PFEM, the fewer iterations are required in the fine-tuning phase. As PFEM is exposed to increasingly diverse geometries, material distributions, and boundary conditions over time, its pretraining model can continuously improve. In this sense, PFEM exhibits a self-learning capability: it becomes progressively more efficient as more scenarios are encountered. This characteristic fundamentally distinguishes PFEM from conventional FEM solvers, which do not possess the ability to improve.
5.6 Comparison between PFEM fine-tuning and preconditioned iterative methods
In the fine-tuning stage of PFEM, the prediction generated by the Transolver model is used as the initial guess for the iterative solver. While PFEM shares certain similarities with classical preconditioned iterative methods, the underlying mechanisms are fundamentally different.
@@ 5.6 Comparison between PFEM fine-tuning and preconditioned iterative methods
The relationship between PFEM fine-tuning and preconditioned iteration can be intuitively understood through a mountain-climbing analogy. Both methods seek to reach the summit, corresponding to the converged solution. A preconditioner reshapes the mountain, reducing steep slopes and making the ascent easier. PFEM, on the other hand, starts the climb from a higher elevation by supplying a better initial guess, thereby shortening the remaining distance to the summit.
@@ 6 Conclusion
In this work, we propose the Pretrain Finite Element Method (PFEM), a novel computational framework that integrates the efficiency of neural operator learning with the accuracy and robustness of classical finite element methods, while eliminating the need for labeled training data. PFEM consists of two stages: a pretraining stage and a fine-tuning stage. In the pretraining stage, a Transolver-based neural operator is trained purely through PDEs, using only point-cloud inputs that encode all essential information for numerical simulation, including geometry (spatial coordinates), material properties (e.g., Young’s modulus and Poisson’s ratio), and boundary conditions. In the fine-tuning stage, the prediction produced by the pretrained operator is used as the initial guess for a conventional iterative solver, thereby accelerating convergence without sacrificing accuracy. We systematically validate PFEM on both linear and nonlinear problems, including elasticity and hyperelasticity benchmarks, under variations in geometry, material properties, and boundary conditions. The numerical results demonstrate that the pretraining stage of PFEM exhibits strong generalization capability across these variations, while the fine-tuning stage significantly reduces the number of iterations required by classical solvers. These results confirm that PFEM effectively combines fast operator-based prediction with high-fidelity numerical refinement.
@@ Appendix B Iterative algorithms of the finite element method
The fine-tuning phase of PFEM then follows the iterative scheme
@@
Appendix C Benefit of the fine-tuning stage in PFEM
Compared with conventional numerical solvers, the fine-tuning stage of PFEM can significantly reduce the number of iterations required for convergence. Below, we provide a convergence analysis for iterative methods applied to linear systems to explain this phenomenon.
@@ Appendix C Benefit of the fine-tuning stage in PFEM
This inequality indicates that, to reach the same target accuracy $\|\boldsymbol{e}^{(k)}\|$, a better initial guess reduces the upper bound of the required iteration number $k$. This provides a theoretical explanation for why a more accurate initial solution from the pretraining stage of PFEM can reduce the iteration count in the fine-tuning stage, especially in scenarios where conventional iterative solvers cannot easily obtain a good initial guess.

### v2

@@ Abstract
We propose a Pretrained Finite Element Method (PFEM), a physics-driven framework that bridges the efficiency of neural operator learning with the accuracy and robustness of classical finite element methods (FEM). PFEM consists of a physics-informed pretraining stage and an optional warm-start stage. In the pretraining stage, a neural operator based on the Transolver architecture is trained solely from governing partial differential equations, without relying on labeled solution data. The model operates directly on unstructured point clouds, jointly encoding geometric information, material properties, and boundary conditions, and produces physically consistent initial solutions with extremely high computational efficiency. PDE constraints are enforced through explicit finite element–based differentiation, avoiding the overhead associated with automatic differentiation. In the warm-start stage, the pretrained prediction is used as an initial guess for conventional FEM solvers, preserving their accuracy, convergence guarantees, and extrapolation capability while substantially reducing the number of iterations required to reach a prescribed tolerance. PFEM is validated on a broad range of benchmark problems, including linear elasticity and nonlinear hyperelasticity with complex geometries, heterogeneous materials, and arbitrary boundary conditions. Numerical results demonstrate strong generalization in the pretraining stage with relative errors on the order of 1%, and speedups of up to one order of magnitude in the warm-start stage compared to FEM with zero initial guesses.
@@
Transolver , Neural operator , Physics-informed neural operator , Physics-informed neural network , AI for PDEs , Solid Mechanics
@@ 5.5 Unlimited generalization and self-learning capability of PFEM
PFEM possesses an inherent capability for unlimited generalization. This property arises from the fact that, in the warm-start stage, PFEM employs conventional FEM-based iterative solvers, which naturally exhibit extrapolation capability beyond the training distribution. Consequently, the performance of PFEM does not fundamentally deteriorate when encountering unseen geometries, material properties, or boundary conditions. In practice, the number of iterations required in the warm-start stage primarily depends on the quality of the initial solution provided by the pretraining stage.
To reduce the computational cost of the warm-start stage, it is therefore crucial that the pretraining phase delivers a high-quality initial guess. Notably, the pretraining phase of PFEM does not require labeled data; it only relies on point-cloud inputs encoding geometry, material properties, and boundary conditions. This observation highlights an important future research direction: how to optimally sample the input space during pretraining?
@@
In summary, the higher the accuracy achieved in the pretraining phase of PFEM, the fewer iterations are required in the warm-start phase. As PFEM is exposed to increasingly diverse geometries, material distributions, and boundary conditions over time, its pretraining model can continuously improve. In this sense, PFEM exhibits a self-learning capability: it becomes progressively more efficient as more scenarios are encountered. This characteristic fundamentally distinguishes PFEM from conventional FEM solvers, which do not possess the ability to improve.
5.6 Comparison between PFEM warm-start and preconditioned iterative methods
In the warm-start stage of PFEM, the prediction generated by the Transolver model is used as the initial guess for the iterative solver. While PFEM shares certain similarities with classical preconditioned iterative methods, the underlying mechanisms are fundamentally different.
@@ 5.6 Comparison between PFEM warm-start and preconditioned iterative methods
The relationship between PFEM warm-start and preconditioned iteration can be intuitively understood through a mountain-climbing analogy. Both methods seek to reach the summit, corresponding to the converged solution. A preconditioner reshapes the mountain, reducing steep slopes and making the ascent easier. PFEM, on the other hand, starts the climb from a higher elevation by supplying a better initial guess, thereby shortening the remaining distance to the summit.
@@ 6 Conclusion
In this work, we propose the Pretrain Finite Element Method (PFEM), a novel computational framework that integrates the efficiency of neural operator learning with the accuracy and robustness of classical finite element methods, while eliminating the need for labeled training data. PFEM consists of two stages: a pretraining stage and a warm-start stage. In the pretraining stage, a Transolver-based neural operator is trained purely through PDEs, using only point-cloud inputs that encode all essential information for numerical simulation, including geometry (spatial coordinates), material properties (e.g., Young’s modulus and Poisson’s ratio), and boundary conditions. In the warm-start stage, the prediction produced by the pretrained operator is used as the initial guess for a conventional iterative solver, thereby accelerating convergence without sacrificing accuracy. We systematically validate PFEM on both linear and nonlinear problems, including elasticity and hyperelasticity benchmarks, under variations in geometry, material properties, and boundary conditions. The numerical results demonstrate that the pretraining stage of PFEM exhibits strong generalization capability across these variations, while the warm-start stage significantly reduces the number of iterations required by classical solvers. These results confirm that PFEM effectively combines fast operator-based prediction with high-fidelity numerical refinement.
@@ Appendix B Iterative algorithms of the finite element method
The warm-start phase of PFEM then follows the iterative scheme
@@
Appendix C Benefit of the warm-start stage in PFEM
Compared with conventional numerical solvers, the warm-start stage of PFEM can significantly reduce the number of iterations required for convergence. Below, we provide a convergence analysis for iterative methods applied to linear systems to explain this phenomenon.
@@ Appendix C Benefit of the warm-start stage in PFEM
This inequality indicates that, to reach the same target accuracy $\|\boldsymbol{e}^{(k)}\|$, a better initial guess reduces the upper bound of the required iteration number $k$. This provides a theoretical explanation for why a more accurate initial solution from the pretraining stage of PFEM can reduce the iteration count in the warm-start stage, especially in scenarios where conventional iterative solvers cannot easily obtain a good initial guess.